# Graph Attention with Random Rewiring

### v1

@@ 1 Introduction
We aim to design a GNN that encompasses the graph inductive bias of message passing, the improved information flow through graph rewiring, and the representation power of the attention mechanism. Towards this end, we propose Graph-Rewiring Attention with Stochastic Structures (GRASS), a novel GNN inspired by these three paradigms. GRASS rewires the input graph by superimposing a random regular graph. At each layer, it applies an attentive node aggregator with attention weights computed from edge representations, and an MLP edge aggregator that updates edge representations to reflect node relationships.
@@ 2 Related Work
The General, Powerful, Scalable (GPS) Graph Transformer [55] represents a hybrid approach, merging MPNNs’ inductive bias with the global perspective of Linear-Time Transformers. Exphormer [60] presents a natural generalization of BigBird to graphs by incorporating local, global, and random attention. It employs random regular expander graphs as an alternative to BigBird’s random attention patterns on sequential data. However, these models, along with prior GTs, focus on adapting Transformer-style dot-product attention to graph-structured data, rather than investigating graph-specialized alternatives. Their performance does not exceed that of more tailored attention mechanisms, such as GRIT [46].
@@
In summary, while existing methodologies have significantly advanced the field of graph neural networks, they are often constrained by limited expressivity within the MPNN framework, computational inefficiencies, or challenges in incorporating graph inductive biases. GRASS emerges as a novel solution that addresses these challenges through an innovative combination of random rewiring and a tailored attention mechanism, offering a competitive solution to graph-structured learning.
@@ 3 Methods
    Permutation Equivariance. Unlike tokens in a sentence or pixels in an image, nodes in a graph are unordered, and therefore the model should be permutation equivariant by construction. Since reordering the nodes of a graph does not change the graph, permuting the nodes of the input graph of a GNN layer should result in the same permutation of its output [65]. Formally, let $f(\mathbf{X},\mathbf{E},\mathbf{A})$ be the function computed by a layer, where $\mathbf{X}:\mathbb{R}^{|V|\times n_{\text{node}}}$ represents node features with $n_{\text{node}}$ dimensions, $\mathbf{E}:\mathbb{R}^{|V|\times|V|\times n_{\text{edge}}}$ represents edge features with $n_{\text{edge}}$ dimensions, and $\mathbf{A}:\{0,1\}^{|V|\times|V|}$ represents the adjacency matrix (edge weights are considered scalar-valued edge features). If the layer is permutation equivariant and $(\mathbf{X}_{\text{out}},\mathbf{E}_{\text{out}})=f(\mathbf{X}_{\text{in}},\mathbf{E}_{\text{in}},\mathbf{A})$, then $(\mathbf{PX}_{\text{out}},\mathbf{PE}_{\text{out}}\mathbf{P}^{\top})=f(\mathbf{PX}_{\text{in}},\mathbf{PE}_{\text{in}}\mathbf{P}^{\top},\mathbf{PAP}^{\top})$ for an arbitrary permutation matrix $\mathbf{P}$.
@@
    Effective Communication – The model should facilitate long-range communication between nodes. Numerous real-world tasks require the GNN to capture interactions between distant nodes [24]. However, MPNN layers, which propagate information locally, frequently fail in this regard [46]. A major challenge is underreaching, where an MPNN with $l$ layers is incapable of supporting communication between two nodes $x_{i},x_{j}$ with distance $\delta(x_{i},x_{j})>l$ [1]. Another challenge is oversquashing, where the structure of the graph forces information from a large set of nodes to squash through a small set of nodes to reach its target [62]. A node with a constant-size feature vector may need to relay information from exponentially many nodes (with respect to model depth), leading to excessive compression of messages in deep MPNNs [1]. Addressing these issues is crucial for effectively capturing long-distance relationships.
@@
    Efficient Computation – The model should have $O(|V|+|E|)$ time complexity for its non-precomputable operations. The number of edges $|E|$ in a connected directed graph with $|V|$ vertices may vary from $|V|$ to $|V|(|V|-1)$. However, in practical scenarios, the number of edges typically falls somewhere between these extremes. GNNs with $O(|V|)$ complexity can be incapable of processing every edge, while those with $O(|V|^{2})$ complexity struggle to scale to large-and-sparse graphs where $|E|\in O(|V|)$. Meanwhile, a model with $O(|V|+|E|)$ complexity is efficient on both dense and sparse graphs. While this is tempting, we notice that computing certain graph properties, such as detecting triangles [69] and finding unweighted all-pair shortest paths [13], cannot be done in $O(|V|+|E|)$ time, which prohibits a model with $O(|V|+|E|)$ time complexity from completing tasks that require these properties. As a compromise, we allow the model to preprocess data without this restriction to capture critical structural properties, as long as the model’s non-precomputable steps remain $O(|V|+|E|)$ time. Precomputable operations only need to be performed once per dataset, regardless of the number of training sessions and epochs, thereby reducing their impact on overall efficiency.
@@
The overall structure of GRASS is illustrated in Figure 2. Prior to training, GRASS precomputes the degree matrix and random walk probabilities of each graph in the dataset. At each training iteration, GRASS randomly rewires the input graph, applies node and edge encodings, and passes the graph through multiple attention layers, producing an output graph with the same structure. If the task is a graph regression or classification task, graph pooling is applied to obtain a single output vector for each graph.
@@ 3.1 Graph Encoding
Degree encoding is another simple yet powerful way to represent structural information [72]. For graphs in which the number of possible degrees is limited, the model could precompute a degree embedding tensor $\mathbf{E}:\mathbb{R}^{D^{+}\times D^{-}\times n}$, where $D^{+}$ and $D^{-}$ denote the maximum out- and in-degree of the dataset, while $\text{d}^{+}(i)$ and $\text{d}^{-}(i)$ represent the out- and in-degree of node $i$.
@@
Before entering attention layers, RRWP encodings are added to both node features and edge features (including edges added by random rewiring), while degree encodings are only added to node features.
@@ 3.2 Random Rewiring
To address the issues of underreaching and oversquashing for Effective Communication, GRASS rewires the input graph by superimposing a random regular graph. While deterministic rewiring and non-regular graphs may also improve connectivity, we present the theoretical advantages of using random regular graphs. We also provide details on random graph generation and input graph rewiring.
@@
The diameter of the input graph upper-bounds the distance between any two nodes, and thus the number of layers for an MPNN to propagate messages between them. Superimposing a random regular graph can drastically decrease the diameter of the input graph. The least integer $d$ that satisfies $(r-1)^{d-1}\geq(2+\varepsilon)r|V|\log|V|$ is the upper bound of the diameter of a random $r$-regular graph with $|V|$ nodes with $r\geq 3$ and $\varepsilon>0$ [7]. Since adding edges to a graph never increases its diameter, the diameter $d$ of the rewired input graph is upper-bounded by $d\in O(\log_{r}|V|)$ when $r\geq 3$. Subsequently, all nodes would be able to communicate with each other with $O(\log_{r}|V|)$ message passing layers, significantly reducing the risk of underreaching even on large graphs. In addition, the diameter of a graph is a trivial upper bound of its effective resistance [26], which has been shown to positively correlate with oversquashing [6]. Intuitively, it upper bounds the “length” of the bottleneck through which messages are passed.
@@
Since oversquashing can be attributed to squeezing too many messages through the fixed-size feature vector of a node [1], increasing the number of internally disjoint paths between two nodes can reduce oversquashing by allowing information to propagate through more nodes in parallel. Intuitively, it increases the “width” of the bottleneck. A random $r$-regular graph with $r\geq 2$ almost certainly has a vertex connectivity of $r$ as $|V|\rightarrow\infty$ [27]. Menger’s Theorem then lower-bounds the number of internally disjoint paths by a graph’s vertex connectivity [30]. This provides another perspective on how the random regular graph reduces oversquashing.
@@
Oversquashing has been shown to decrease as the spectral gap of a graph increases, which is defined as $\lambda_{1}$, the smallest positive eigenvalue of the graph’s Laplacian matrix [37]. It has been proven that a random $r$-regular graph sampled uniformly from the set of all $r$-regular graphs with $|V|$ nodes almost certainly has $\mu<2\sqrt{r-1}+1$ as $|V|\rightarrow\infty$, where $\mu$ is the largest absolute value of nontrivial eigenvalues of its adjacency matrix [52]. Since the graph is $r$-regular, its $i$-th adjacency matrix eigenvalue $\mu_{i}$ and $i$-th Laplacian matrix eigenvalue $\lambda_{i}$ satisfy $\lambda_{i}=r-\mu_{i}$ [45], lower-bounding the spectral gap with $\lambda_{1}>r-2\sqrt{r-1}-1$. Therefore, the random regular graph alleviates oversquashing on large input graphs by increasing the spectral gap of the rewired graph.
@@
where $\sqcup$ denotes the disjoint union of sets. The resulting graph $\tilde{R}$ is a random regular pseudograph, and the probability that $\tilde{R}$ is any regular pseudograph with $|V_{G}|$ nodes and degree $r$ is equal [7]. Being a pseudograph, $\tilde{R}$ might not be simple: it might have self-loops and multi-edges. Even when $|V_{G}|$ is large, the probability that $\tilde{R}$ is simple—that it does not have multiple-edges and self-loops—would not be prohibitively small. In particular, it has an asymptotically tight lower bound [7]:
@@
Therefore, if we re-generate $\tilde{R}$ when it’s not simple, the expected number of trials required for us to obtain a simple $\tilde{R}$ is upper-bounded by $e^{\frac{r^{2}}{2}+r}$, which can be kept low by keeping $r$ low. In practice, GRASS would like the generated graph to be simple to avoid unwanted redundancy in the attention mechanism. In particular, self-loops result in a duplicate path of message passing from a node to itself, and multiple-edges result in identical activation on these edges. However, the regularity of the graph is desired but not strictly required. Therefore, when $\tilde{R}$ is not simple, we simply remove self-loops and multi-edges from $\tilde{R}$ to obtain $R$, which is always simple but not necessarily regular.
@@
The added edges $E_{R}$ are given a distinct embedding to make them distinguishable from the existing edges $E_{G}$. This aids Selective Aggregation, as it enables a node to select between its neighbors and a random node. Edge encodings are also given to the added edges, which is instrumental in Relationship Representation. Although an added edge $(i,j)\in E_{R}$ lacks edge features that represent semantic relationships in the input graph, the structural relationship between nodes $i$ and $j$ is represented by the random walk probabilities $\mathbf{P}_{:,i,j}$ given to that edge as its RRWP encoding.
@@ 3.3 Attention Mechanism
In calculating attention weights, GRASS scales the input to the softmax operation by the logarithm of the number of keys, which is equal to the in-degree of that node after the random edges are added, as it has been found to help the attention mechanism focus when there is a large number of keys by reducing the influence of the number of keys on the entropy of the attention weight distribution [17], which benefits Selective Aggregation. Intuitively, this allows a node to focus its attention on relevant in-neighbors regardless of the number of in-neighbors it has. Since the number of keys a node attends to is equal to its in-degree, this approach is also analogous to degree scaling [19], which is widely used in MPNNs to incorporate degree information into the attention mechanism.
In addition, GRASS uses DropKey [41] as a regularizer, which, in the context of graph attention, is effective in reducing dependence on particular edges that appear in the graph. This improves the robustness against structural noise and thus facilitates Selective Aggregation. DropKey can be seen as the random removal of edges because it randomly masks the attention matrix prior to the softmax operation. Thus, it elegantly complements random rewiring, which only adds edges. This equivalence could not be established for attention dropout [75] despite its popularity in Transformers.
Like Transformers, the output of the attention mechanism is passed through an FFN. $\phi$ denotes a ReLU-like [49, 54, 48] nonlinear activation function.
@@
PN-V is a variant of Power Normalization, which normalizes the root mean square of the batch and is more stable than BN on large-variance inputs [58]. BN has been shown to work poorly in GNNs [12], which would be made worse by the variance introduced by random rewiring. PN-V is used by GRASS to address this problem, allowing it to achieve adequate stability during training while still enjoying normalized batch distributions [58].
The residual connection [33] is scaled by $\alpha$, a coefficient calculated according to DeepNorm [67], which stabilizes the training of deep residual networks. This allows us to use post-normalization, which has been shown to increase expressivity while avoiding the problem of exploding model update [44, 67]. This is essential for us to validate the robustness of GRASS against oversquashing and oversmoothing, which deep models are vulnerable to, by training the GNN with around 50 layers.
@@
While the proposed attention mechanism naturally achieves Directionality Preservation by aggregating information in the same direction as edges, it can severely restrict the flow of information, putting it in conflict with Effective Communication. As a solution, the directionality of all edges is alternated across layers: their directions are identical to those of the rewired graph $H$ in odd layers and inverted in even layers. This enables the model to propagate information in both directions when the input graph is directed, which improves the model’s expressivity [15]. We empirically find this to be beneficial on undirected graphs as well.

### v2

@@ 1 Introduction
We aim to design a GNN that encompasses the graph inductive bias offered by message passing, the improved information flow achieved through graph rewiring, and the representation power of the attention mechanism. Towards this end, we propose Graph-Rewiring Attention with Stochastic Structures (GRASS), a novel GNN inspired by these three paradigms. GRASS rewires the input graph by superimposing a random regular graph. At each layer, it applies an attentive node aggregator with attention weights computed from edge representations, and an MLP edge aggregator that updates edge representations to reflect node relationships.
@@ 2 Related Work
The General, Powerful, Scalable (GPS) Graph Transformer [56] represents a hybrid approach, merging MPNNs’ inductive bias with the global perspective of Linear-Time Transformers. Exphormer [61] presents a natural generalization of BigBird to graphs by incorporating local, global, and random attention. It employs random regular expander graphs as an alternative to BigBird’s random attention patterns on sequential data, and is analogous to the random rewiring mechanism of GRASS. However, these models, along with prior GTs, focus on adapting Transformer-style dot-product attention to graph-structured data, rather than investigating graph-specialized alternatives. Their performance does not exceed that of more tailored attention mechanisms, such as GRIT [47], on a variety of tasks.
@@
In summary, while existing methodologies have significantly advanced the field of graph neural networks, they are often constrained by limited expressivity within the MPNN framework, computational inefficiencies, or challenges in incorporating a graph inductive bias. GRASS emerges as a novel solution that addresses these challenges through the combination of random rewiring and a graph-tailored attention mechanism, offering a competitive solution to graph-structured learning.
@@ 3 Methods
    Permutation Equivariance. Unlike tokens in a sentence or pixels in an image, nodes in a graph are unordered, and therefore the model should be permutation equivariant by construction. Since reordering the nodes of a graph does not change the graph, permuting the nodes of the input graph of a GNN layer should only result in the same permutation of its output [67]. Formally, let $f(\mathbf{X},\mathbf{E},\mathbf{A})$ be the function computed by a layer, where $\mathbf{X}:\mathbb{R}^{|V|\times n_{\text{node}}}$ represents node features with $n_{\text{node}}$ dimensions, $\mathbf{E}:\mathbb{R}^{|V|\times|V|\times n_{\text{edge}}}$ represents edge features with $n_{\text{edge}}$ dimensions, and $\mathbf{A}:\{0,1\}^{|V|\times|V|}$ represents the adjacency matrix (edge weights are considered scalar-valued edge features). If the layer is permutation equivariant and $(\mathbf{X}_{\text{out}},\mathbf{E}_{\text{out}})=f(\mathbf{X}_{\text{in}},\mathbf{E}_{\text{in}},\mathbf{A})$, then $(\mathbf{PX}_{\text{out}},\mathbf{PE}_{\text{out}}\mathbf{P}^{\top})=f(\mathbf{PX}_{\text{in}},\mathbf{PE}_{\text{in}}\mathbf{P}^{\top},\mathbf{PAP}^{\top})$ for an arbitrary permutation matrix $\mathbf{P}$.
@@
    Effective Communication – The model should facilitate long-range communication between nodes. Numerous real-world tasks require the GNN to capture interactions between distant nodes [24]. However, MPNN layers, which propagate information locally, frequently fail in this regard [47]. A major challenge is underreaching, where an MPNN with $l$ layers is incapable of supporting communication between two nodes $x_{i},x_{j}$ with distance $\delta(x_{i},x_{j})>l$ [1]. Another challenge is oversquashing, where the structure of the graph forces information from a large set of nodes to squeeze through a small set of nodes to reach its target [64]. A node with a constant-size feature vector may need to relay information from exponentially many nodes (with respect to model depth), leading to excessive compression of messages in deep MPNNs [1]. Addressing these issues is crucial for effectively capturing long-distance relationships.
@@
    Efficient Computation – The model should have $O(|V|+|E|)$ time complexity for its non-precomputable operations. The number of edges $|E|$ in a connected directed graph with $|V|$ vertices may vary from $|V|$ to $|V|(|V|-1)$. However, in practical scenarios, the number of edges typically falls somewhere between these extremes. GNNs with $O(|V|)$ complexity can be incapable of processing every edge, while those with $O(|V|^{2})$ complexity struggle to scale to large-and-sparse graphs where $|E|\in O(|V|)$. Meanwhile, a model with $O(|V|+|E|)$ complexity can be efficient on both dense and sparse graphs. While this is tempting, we notice that computing certain graph properties, such as detecting triangles [71] and finding unweighted all-pair shortest paths [14], cannot be done in $O(|V|+|E|)$ time, which prohibits a model with $O(|V|+|E|)$ time complexity from completing tasks that require these properties. As a compromise, we allow the model to preprocess data without this restriction to capture critical structural properties, as long as the model’s non-precomputable steps remain $O(|V|+|E|)$ time. Precomputable operations only need to be performed once per dataset, regardless of the number of training sessions and epochs, thereby reducing their impact on overall efficiency.
@@
The high-level structure of GRASS is illustrated in Figure 2. Prior to training, GRASS precomputes the degree matrix and random walk probabilities of each graph in the dataset. At each training iteration, GRASS randomly rewires the input graph, applies node and edge encodings, and passes the graph through multiple attention layers, producing an output graph with the same structure. For tasks that require a graph-level representation, such as graph regression and graph classification, pooling is performed on the output graph to obtain a single output vector.
@@ 3.1 Graph Encoding
Degree encoding is another simple yet powerful way to represent structural information [74]. For datasets in which node degrees can take a small number of values, the model could learn a degree embedding tensor $\mathbf{E}:\mathbb{R}^{D^{+}\times D^{-}\times n}$, where $D^{+}$ and $D^{-}$ denote the maximum out- and in-degree of the dataset, while $\text{d}^{+}(i)$ and $\text{d}^{-}(i)$ represent the out- and in-degree of node $i$.
@@
Before entering attention layers, RRWP encodings are added to both node features and edge features (including those of edges added by random rewiring), while degree encodings are only added to node features.
@@ 3.2 Random Rewiring
To address the issues of underreaching and oversquashing for Effective Communication, GRASS rewires the input graph by superimposing a random regular graph. While deterministic rewiring and non-regular graphs may also improve connectivity, we present potential advantages of using random regular graphs. We also provide details on random regular graph generation and input graph rewiring.
@@
The diameter of a graph upper-bounds the distance between two nodes, and thus the number of layers for an MPNN to propagate information between them [1]. Superimposing a random regular graph onto the input graph can drastically decrease its diameter. The least integer $d$ that satisfies $(r-1)^{d-1}\geq(2+\varepsilon)r|V|\log|V|$ is the upper bound of the diameter of almost every random $r$-regular graph with $|V|$ nodes, where $r\geq 3$ and $\varepsilon>0$ [8]. Since adding edges to a graph never increases its diameter, the diameter $d$ of the rewired graph is asymptotically upper-bounded by $d\in O(\log_{r}|V|)$ when $r\geq 3$. Subsequently, all nodes would be able to communicate with each other given $O(\log_{r}|V|)$ message passing layers, which could significantly reduce the risk of underreaching on large graphs. In addition, the diameter of a graph is a trivial upper bound of its effective resistance [26], which has been shown to be positively associated with oversquashing [7]. Intuitively, it upper bounds the “length” of the bottleneck through which messages are passed.
@@
Since oversquashing can be attributed to squeezing too many messages through the fixed-size feature vector of a node [1], increasing the number of internally disjoint paths between two nodes may reduce oversquashing by allowing information to propagate through more nodes in parallel. Intuitively, it increases the “width” of the bottleneck. A random $r$-regular graph with $r\geq 2$ almost certainly has a vertex connectivity of $r$ as $|V|\rightarrow\infty$ [27]. Menger’s Theorem then lower-bounds the number of internally disjoint paths by a graph’s vertex connectivity [31]. This provides another perspective on how the random regular graph may reduce oversquashing.
@@
Oversquashing has been shown to decrease as the spectral gap of a graph increases, which is defined as $\lambda_{1}$, the smallest positive eigenvalue of the graph’s Laplacian matrix [38]. It has been proven that a random $r$-regular graph sampled uniformly from the set of all $r$-regular graphs with $|V|$ nodes almost certainly has $\mu<2\sqrt{r-1}+1$ as $|V|\rightarrow\infty$, where $\mu$ is the largest absolute value of nontrivial eigenvalues of its adjacency matrix [53]. Since the graph is $r$-regular, its $i$-th adjacency matrix eigenvalue $\mu_{i}$ and $i$-th Laplacian matrix eigenvalue $\lambda_{i}$ satisfy $\lambda_{i}=r-\mu_{i}$ [46], lower-bounding the spectral gap with $\lambda_{1}>r-2\sqrt{r-1}-1$. Therefore, the superimposed random regular graph may alleviate oversquashing by increasing the spectral gap of the rewired graph.
@@
Here, $\sqcup$ denotes the disjoint union of sets. The resulting graph $\tilde{R}$ is a random regular pseudograph, and the probability that $\tilde{R}$ is any regular pseudograph with $|V_{G}|$ nodes and degree $r$ is equal [8]. Being a pseudograph, $\tilde{R}$ might not be simple: it might contain self-loops and multi-edges. Even when $|V_{G}|$ is large, the probability that $\tilde{R}$ is simple—that it does not contain self-loops or multi-edges—would not be prohibitively small. In particular, it has an asymptotically tight lower bound [8]:
@@
Therefore, if we re-generate $\tilde{R}$ when it’s not simple, the expected number of trials required for us to obtain a simple $\tilde{R}$ is upper-bounded by $e^{\frac{r^{2}}{2}+r}$, which can be kept low by keeping $r$ low. In practice, GRASS would like the superimposed graph to be simple, in order to avoid unwanted redundancy in the attention mechanism. In particular, self-loops result in a duplicate path of message passing from a node to itself, while multi-edges result in multiple identical messages. Meanwhile, the regularity of the graph is desired but not strictly required. Therefore, when $\tilde{R}$ is not simple, we simply remove self-loops and multi-edges from $\tilde{R}$ to obtain $R$, which is always simple but not necessarily regular.
@@
The added edges $E_{R}$ are given a distinct embedding to make them distinguishable from the existing edges $E_{G}$. This aids Selective Aggregation, as it enables a node to select between its neighbors and a random node. Edge encodings are also given to the added edges, which are instrumental in Relationship Representation. Although an added edge $(i,j)\in E_{R}$ lacks edge features that represent semantic relationships in the input graph, the structural relationship between nodes $i$ and $j$ is represented by the random walk probabilities $\mathbf{P}_{:,i,j}$ given to that edge as its RRWP encoding.
@@ 3.3 Attention Mechanism
In calculating attention weights, GRASS scales the input to the softmax operation by the logarithm of the number of keys at each node, which is equal to the in-degree of that node after random rewiring. This has been shown to reduce the impact of the number of keys on the entropy of the distribution of attention weights when there is a large number of keys [17], which supports Selective Aggregation. Intuitively, this allows a node to concentrate its attention weight on relevant in-neighbors regardless of the number of in-neighbors it has. Since the number of keys a node attends to is equal to its in-degree, this approach is also analogous to degree scaling [19], which is widely used in MPNNs to incorporate degree information into message passing.
In addition, GRASS uses DropKey [42] as a regularizer, which, in the context of graph attention, is effective in reducing its dependence on the presence of particular edges in the graph. This improves the robustness of GRASS against structural noise and thus facilitates Selective Aggregation. DropKey can be seen as the random removal of edges because it randomly masks the attention matrix prior to the softmax operation. Thus, it elegantly complements random rewiring, which only adds edges. This equivalence could not be established for Attention Dropout [77] despite its popularity in Transformers.
Like Transformers, the output of the attention mechanism is passed through an FFN. Here, $\phi$ denotes a ReLU-like [50, 55, 49] nonlinear activation function.
@@
PN-V is a variant of Power Normalization [59], which normalizes the root mean square of the batch and is more stable than BN on large-variance inputs. BN has been shown to work poorly in GNNs [13], which would be made worse by the variance introduced with random rewiring. PN-V is used by GRASS to address this problem, allowing it to achieve adequate stability during training while still enjoying normalized batch distributions [59].
The residual connection [34] is scaled by $\alpha$, a coefficient calculated according to DeepNorm [69], which stabilizes the training of deep residual networks. This allows us to use post-normalization, which has been shown to increase expressivity while avoiding the problem of exploding model update [45, 69]. This is essential for us to validate the robustness of GRASS against oversquashing and oversmoothing, which deep GNNs are more vulnerable to, by allowing us to train a deeper model.
@@
While the proposed attention mechanism naturally achieves Directionality Preservation by aggregating information in the same direction as edges, it can severely restrict the flow of information, putting it in conflict with Effective Communication. As a solution, the direction of each edge is switched from one layer to the next: in odd-numbered layers, the directions match those of the edges in the rewired graph $H$, whereas in even-numbered layers, they are reversed. This enables the model to propagate information in both directions even when the input graph is directed, improving its expressivity [58]. We empirically find edge flipping to be beneficial on undirected graphs (where an undirected edge is represented as a pair of directed edges with opposite directions) as well.